# ScaleRAFT: Cross-Scale Recurrent All-Pairs Field Transforms for 3D Motion Estimation

### v1

@@ III-A Feature Extraction
Feature Encoder. We use a convolutional network $g_{\theta}$ with 6 residual connections to extract features. As the number of layers increases, the feature channel gradually becomes wider, and the image size gradually becomes smaller, where $g_{\theta}:\mathbb{R}^{H\times W\times 3}\rightarrow\mathbb{R}^{(H/8)\times(W/8)\times D}(D=256)$.
Context Encoder. Following RAFT, we build a context encoder $h_{\theta}(D=384)$ with the same structure as $g_{\theta}$. This context encoder extracts context features $F_{context}(D=192)$ and implicit features $h_{0}(D=192)$ from $I_{1}$, where the implicit features are used to initialize the internal parameters of GRU optimizer.
III-B Cross-scale Correlation Volume (CSCV)
@@
The Input of CSCV. The main function of CSCV is to sample the correlation features of the current input 3D flow field. Specifically, the input 3D flow field is composed of an optical flow field $(f_{1},f_{2})$ and a scale change field $f_{3}$. The optical flow field describes the displacement of pixels between two frames, and the scale change field is the reciprocal of optical expansion (OE), which describes the pixel scaling factor of the second frame for perfect matching the first frame.
@@
where $I_{2}^{s}\in\mathbb{R}^{(H\times s)\times(W\times s)\times 3}$, $S$ specifies the degree of refinement of the input scale, which is generally set to 4 in our experiment.
Multi Scale Correlation Volume. The multi scale correlation volume $C_{m}$ is a collection of multiple 4D correlation matrices. The construction of the 4D matrix follows RAFT[20], which can be sampled by the coordinates of a pair of matching points. The sample result is the correlation of these two points. Specifically, for the given image features $g_{\theta}(I_{2}^{s})\in\mathbb{R}^{(H\times s)\times(W\times s)\times D}$ and $g_{\theta}(I_{1})\in\mathbb{R}^{H\times W\times D}$ , dot them to obtain a single 4D correlation matrix $c_{s}$:
$\begin{gathered}c_{s}\left(g_{\theta}\left(I_{2}^{s}\right),g_{\theta}\left(I_{1}\right)\right)\in\mathbb{R}^{H\times W\times(H\times s)\times(W\times s)}\\ (c_{s})_{ijkl}=\sum_{h}g_{\theta}\left(I_{2}^{s}\right)_{ijh}\cdot g_{\theta}\left(I_{1}\right)_{klh}\end{gathered}$ (2)
@@
Optical Flow Lookup Operator. We define the optical flow lookup operator $L_{c}^{of}$ to sample the correlation features corresponding to the current optical flow field $(f_{1},f_{2})$ from $C_{m}$.
@@
where $cf_{s}\in\mathbb{R}^{H\times W\times(2r_{f}+1)\times(2r_{f}+1)}$.
@@
Scale Lookup Operator. We define the scale lookup operator $L_{c}^{s}$ to sample the correlation features corresponding to the current scale change field $f_{3}$ from $F_{multi}$.
@@
Based on the sampling coordinates provided by $N_{x^{\prime}}^{s}$, we sample in $F_{multi}$ based on bilinear interpolation, and splice the results into scale correlation feature $F_{scale}\in\mathbb{R}^{H\times W\times(2r_{f}+1)\times(2r_{f}+1)\times(2r_{s}+1)}$.
We follow the RAFT strategy to extract the optical flow correlation feature $F_{of}\in\mathbb{R}^{H\times W\times(2r_{f}+1)\times(2r_{f}+1)\times 4}$, which is sampled from $c_{1}$ ( $2\times 2$ pooling for three times) by $L_{c}^{of}$.
III-C Updating Displacement Field
Initialization. We initialize the optical flow field $(f_{1}^{0},f_{2}^{0})$ to 0 everywhere and scale change field $f_{3}^{0}$ to 1 everywhere, initial $h_{0}$ is given by context encoder.
Updating. We use the residual update to estimate the displacement field $f^{k+1}=f^{k}+\Delta f$, which is based on the gating activation unit GRU, and replace the full connection layer with the convolution layer:
$\mathrm{z}_{\mathrm{t}}=Sigmod\left(Conv\left(W_{z}\cdot\left[h_{t-1},x_{t}\right]\right)\right)$ (7)
$\mathrm{r}_{\mathrm{t}}=Sigmod\left(Conv\left(W_{r}\cdot\left[h_{t-1},x_{t}\right]\right)\right)$
$\tilde{h}_{t}=\tanh\left(Conv\left(W\cdot\left[r_{t}*h_{t-1},x_{t}\right]\right)\right)$
$h_{t}=\left(1-z_{t}\right)*h_{t-1}+z_{t}*\tilde{h}_{t}$
Eq (7) shows the process of the GRU optimizer iteratively updating the hidden parameter $h_{t}$, where the input is $h_{t-1}$ and $x_{t}$, the output is the updated $h_{t}$, and $W_{z},W_{r},W$ is the parameter to be learned. $x_{t}$ is the combination of displacement field $f=(f_{1},f_{2},f_{3})$, correlation feature $F_{of}$ and $F_{scale}$, and context feature $F_{context}$.
The residual $\Delta f=(\Delta f_{1},\Delta f_{2},\Delta f_{3})$ for updating is estimated from $h_{t}$ by a simple two-level convolution module:
$\begin{split}&(\Delta f_{1},\Delta f_{2})=Conv(ReLU(Conv(h_{t})))\\ &\Delta f_{3}=\tanh(Conv(ReLU(Conv(h_{t}))))\end{split}$ (8)
Considering that the scale field requires more detailed estimation, we replaced the GRU with a small Unet[78] network and performed an additional update of $f_{3}$.
Upsampling. The network prediction output is $1/8$ resolution. The results used for prediction and training need upsampling. Here we follow the previous work[20], linearly weight upsampling based on the $3\times 3$ neighborhood of coarse-resolution pixels.
III-D Hybrid Training Pipeline
@@
Note that the captured image blocks here lean towards random textures, as we found that using foreground flying objects with clear semantic information (such as complete car, human, and animal contours) does not perform as well as random textures. Specifically, when occlusion occurs, the network becomes overly confident in inferring the motion of the occluded part of the object as planar motion. Random textures can encourage networks to calculate the motion of these flying foreground based on texture matching, and balance the learning weight with occlusion regression.
Calculate Foreground Labels. Before calculating the label for the flight foreground, it is necessary first to determine whether there is an occlusion in the foreground. Because the flight foreground is primarily a random texture, the network cannot make reasonable inferences about the occluded parts. We propose a difference rate $Df$ indicator to measure whether the prospect is retained:
@@
where $N_{q}$ and $N_{r}$ represent the number of foreground flying object pixels in the first and second frames, respectively. In the experiment, we only used flight prospects with $Df$ less than 0.3 to ensure no significant obstructions or great scale changes during the flight.
@@
Optical Flow Loss. We supervise the optical flow results with $L_{1}$ distance between prediction and ground truth, the optical loss $\mathcal{L}_{f}$ is defined as:
@@
we set $\gamma=0.8$, $N=12$ in our experiments, the $\bm{f_{gt}}=(f_{1gt},f_{2gt})$ is composed of a hybrid of truth labels from the dataset and the generated $\bm{f_{fg}}$ labels.
@@
where $\bm{\tau}_{gt}$ is also a hybrid of true values $\bm{\tau}$ and generated labels $\bm{\tau}_{fg}$, $\bm{\tau}=Z^{\prime}_{gt}/Z_{gt}$, $Z^{\prime}_{gt}$ and $Z_{gt}$ are the ground truth depths of matching pixels in the second frame and the first frame, we set $\gamma=0.8$, $N=13$ in our experiments.
@@
III-E Convert ScaleRAFT output to 3D motion
Referring to previous work[18], we converted the 3D motion $(f_{1},f_{2},f_{3})$ into motion-in-depth and scene flow. Firstly, based on Yang’s[18] derivation, we can obtain the following approximate equation:
@@ III-E Convert ScaleRAFT output to 3D motion
where $\bm{Cd}$ is the image coordinate index matrix, normalize scene flow $\overline{\bm{t}}=\bm{K^{-1}}[(f_{3}(\bm{p})-1)\bm{p}+f_{3}(\bm{p})\bm{f}]$. We have presented a detailed derivation process in the Supplementary Material. I.
III-F Expand ScaleRAFT to RGB-D Method
In this section, we used a simple yet practical rigid background optimization (RBO) approach to upgrade ScaleRAFT to the RGBD version.
Specifically, we first remove manually specified motion foreground objects (cars and pedestrians in KITTI) from the current frame based on semantic masks [79] to obtain the background point set $\bm{p}_{2d}^{1}$. Then, based on depth values $Z$ and optical flow $\bm{f}=(f_{1},f_{2})$, we calculate the 3D coordinates $\bm{P}_{3d}^{1}$ of pixels in the first frame and their corresponding planar coordinates $\bm{p}_{2d}^{2}$ in the second frame:
$\bm{P}_{3d}^{1}=Z\bm{K^{-1}}\bm{p}_{2d}^{1}$ (21)
$\bm{p}_{2d}^{2}=\bm{p}_{2d}^{1}+\bm{f}(\bm{p}_{2d}^{1})$ (22)
At this point, the rigid background optimization problem is transformed into a PnP problem. We use the $solvePnP$ function in OpenCV [80, 81] to solve the corresponding rotation and translation between two frames:
$\bm{R}_{bg},\bm{T}_{bg}=solvePnP(\bm{P}_{3d}^{1},\bm{p}_{2d}^{2},\bm{K})$ (23)
Based on $\bm{R}_{bg}$ and $\bm{T}_{bg}$, we consider the background pixels as a whole and recalculate the 3D motion $\bm{t}_{bg}$ of the background:
$\bm{P}_{3d}^{2}=\bm{R}_{bg}\bm{P}_{3d}^{1}+\bm{T}_{bg}$ (24)
$\bm{t}_{bg}=\bm{P}_{3d}^{2}-\bm{P}_{3d}^{1}$ (25)
where $\bm{t}_{bg}$ is the optimized background scene flow, and we directly replace the original background result with it. The RBO idea has been indirectly or directly used in most current scene flow methods. Our RBO algorithm refers to the background optimization section in [14]. More details about RBO can be found in [13, 15, 14].
Why RBO? The main advantage of the cross-scale matching algorithm ScaleRAFT is that it can better estimate the moving foreground, while the most significant advantage of rigid optimization technology is that it can optimize the entire static background as a rigid object at once. These two algorithms are complementary in terms of characteristics. The outstanding performance demonstrated by ScaleRAFT-RBO in Tab. II demonstrates the effectiveness of this complementary combination.
III-G Apply CSCV to RAFT3D
RAFT3D[13] is a state-of-the-art method for scene flow. It regresses scene flows directly from RGBD inputs based on RAFT baselines and iteratively optimizes the regression results through a rigid optimization layer. We extend CSCV to RAFT3D baseline (called ScaleRAFT3D) to further validate the effectiveness of cross-scale matching idea.
The overall structure of ScaleRAFT3D is similar to ScaleRAFT, both of which use CSCV to extract cross-scale matching features and iteratively optimize the 3D motion. The main difference is that ScaleRAFT3D adds a rigid optimization layer after the GRU optimizer and directly estimates the $SE(3)$ transformation field $\mathbf{T}$. In addition, in order to extract semantic information for rigid optimization, the original 6-layer convolutional context extractor has been replaced with Resnet50[82]. We have presented a detailed ScaleRAFT3D structure in the Supplementary Material Sec. II.
@@ IV Experiments
We train and evaluate our method on the standard scene flow datasets including KITTI, Driving, Flyingthings3D, and MPI-Sintel. In section IV-A, we will detail our benchmark datasets and training settings. In section IV-B, we evaluate the performance of ScaleRAFT on the MID task, which is also the core indicator of monocular 3D flow. In section IV-C, we evaluate ScaleRAFT on the TTC task. In section IV-D, we submit the proposed methods to the KITTI’s scene flow benchmark [^3] http://www.cvlibs.net/datasets/kitti/eval_scene_flow.php for comparison with the state-of-the-art scene flow methods. In section IV-E, we tested the performance of ScaleRAFT and ScaleRAFT3D on the Flyingthings3D. Finally, in section IV-F, we conduct ablation experiments.

### v2

@@ III-A Feature Extraction
We replaced the residual network in RAFT with ResNet18[54], which has the advantage of being able to use pre-trained weights from MAE[55] or ImageNet[56]. The results indicate that this can improve training efficiency and introduce additional prior knowledge to improve generalization performance. Specifically, we constructed a feature encoder and a context encoder, respectively:
Feature Encoder. For the feature encoder $g_{\theta}$, we take its truncated outputs at 1/8 and 1/16 scales, where $g_{\theta}^{1}:\mathbb{R}^{H\times W\times 3}\rightarrow\mathbb{R}^{(H/8)\times(W/8)\times D}$ and $g_{\theta}^{2}:\mathbb{R}^{H\times W\times 3}\rightarrow\mathbb{R}^{(H/16)\times(W/16)\times D}(D=256)$.
Context Encoder. Following RAFT, we build a context encoder $h_{\theta}$, whose input is the superposition of frames $I_{1}$ and $I_{2}$, and the rest of the structure is the same as $g_{\theta}$. The output of the 1/16 scale $F_{init}(D=384)$ is directly used to initialize, while the output of the 1/8 scale is divided into two parts: one is the context feature $F_{context}(D=192)$ and the other is the implicit feature $h_{0}(D=192)$ used for GIR iteration.
III-B Initialization of Optical Flow and Scale Flow
RAFT-type methods usually initialize the flow field to 0 and then refine it iteratively through a GRU module. However, we found that this leads to two problems. Firstly, GRU iterative networks must simultaneously complete tasks such as large field optimization (early iteration) and fine-tuning (late iteration), significantly increasing the difficulty of training the iterative network. Secondly, due to limitations in memory and cost, it is not easy to increase the sampling range of the correlation volume during iteration in order to obtain a larger receptive field.
In ScaleFlow++, we drew on the successful experiences of IGEV[57] and SEA-RAFT[58] and used an independent lightweight network to quickly initialize the 3D flow, reducing the number of iterations of the optimizer and alleviating task differences at different iteration stages. Specifically, the input for initializing the network is the contextual feature $F_{init}$ and the correlation feature $F_{of}^{init}$. We first use the GIR module proposed in Sec. III-D to encode the above features:
$h_{init}=\textbf{{GIR}}(F_{of}^{init},F_{init})$ (1)
where $F_{of}^{init}\in\mathbb{R}^{H\times W\times(2r+1)\times(2r+1)}$ is sampled from a 1/16 scale 4D correlation volume, the sampling optical flow field is all 0, and the sampling radius $r$ is 6.
Next, we upsample the hi to 1/8 scale and use a simple two-layer convolution head to predict the initialized optical flow field $(f_{1}^{0},f_{2}^{0})$ and scale change field $f_{3}^{0}$:
$\begin{split}(f_{1}^{0},f_{2}^{0})&=\textrm{Conv}(\textrm{ReLU}(\textrm{Conv}(\textrm{Upsample}(h_{init}))))\\ f_{3}^{0}&=\textrm{Conv}(\textrm{ReLU}(\textrm{Conv}(\textrm{Upsample}(h_{init}))))\end{split}$ (2)
III-C Cross-scale Correlation Volume (CSCV)
@@
III-C1 The Input of CSCV
The main function of CSCV is to sample the correlation features of the current input 3D flow field. Specifically, the input 3D flow field is composed of an optical flow field $(f_{1},f_{2})$ and a scale change field $f_{3}$. The optical flow field describes the displacement of pixels between two frames, and the scale change field is the reciprocal of optical expansion (OE), which describes the pixel scaling factor of the second frame for perfect matching the first frame.
@@ III-C1 The Input of CSCV
where $I_{2}^{s}\in\mathbb{R}^{(8\times H\times s)\times(8\times W\times s)\times 3}$, $S$ specifies the degree of refinement of the input scale, which is generally set to 4 in our experiment. Because subsequent iterations were conducted on the 1/8 scale, the width and height were multiplied by a coefficient of 8 to facilitate the expression of following functions.
Figure 5: Global Iterative Refinement Module (GIR). Similar to Unet, GIR is mainly composed of funnel-shaped encoders and decoders. Specifically, GIR uses ConNextV2 with large kernel convolution characteristics at multiple scales, where a single convolution kernel can cover almost the entire frame. Combined with the pyramid pooling module at the smallest scale, it greatly enhances global perception capability.
III-C2 Multi Scale Correlation Volume
The multi scale correlation volume $C_{m}$ is a collection of multiple 4D correlation matrices. The construction of the 4D matrix follows RAFT, which can be sampled by the coordinates of a pair of matching points. The sample result is the correlation of these two points. Specifically, for the given image features $g_{\theta}^{1}(I_{2}^{s})\in\mathbb{R}^{(H\times s)\times(W\times s)\times D}$ and $g_{\theta}^{1}(I_{1})\in\mathbb{R}^{H\times W\times D}$ , dot them to obtain a single 4D correlation matrix $c_{s}$:
$\begin{gathered}c_{s}\left(g_{\theta}^{1}\left(I_{2}^{s}\right),g_{\theta}^{1}\left(I_{1}\right)\right)\in\mathbb{R}^{H\times W\times(H\times s)\times(W\times s)}\\ (c_{s})_{ijkl}=\sum_{h}g_{\theta}^{1}\left(I_{2}^{s}\right)_{ijh}\cdot g_{\theta}^{1}\left(I_{1}\right)_{klh}\end{gathered}$ (4)
@@ III-C2 Multi Scale Correlation Volume
III-C3 Optical Flow Lookup Operator
We define the optical flow lookup operator $L_{c}^{of}$ to sample the correlation features corresponding to the current optical flow field $(f_{1},f_{2})$ from $C_{m}$.
@@ III-C3 Optical Flow Lookup Operator
where $cf_{s}\in\mathbb{R}^{H\times W\times(2r_{f}+1)\times(2r_{f}+1)}$. In Fig. 6, we demonstrate the working principle of $F_{of}$ through a simplified cross-scale matching example.
@@
III-C4 Scale Lookup Operator.
We define the scale lookup operator $L_{c}^{s}$ to sample the correlation features corresponding to the current scale change field $f_{3}$ from $F_{multi}$.
@@
Based on the sampling coordinates provided by $N_{x^{\prime}}^{s}$, we sample in $F_{multi}$ based on bilinear interpolation, and splice the results into scale correlation feature $F_{scale}\in\mathbb{R}^{H\times W\times(2r_{f}+1)\times(2r_{f}+1)\times(2r_{s}+1)}$. Moreover, we also follow the RAFT strategy to extract the optical flow correlation feature $F_{of}\in\mathbb{R}^{H\times W\times(2r_{f}+1)\times(2r_{f}+1)\times 4}$, which is sampled from $c_{1}$ ( $2\times 2$ pooling for three times) by $L_{c}^{of}$.
III-D Global Iterative Refinement
III-D1 Optimizer
In most RAFT-type methods, the optimizer’s field of view is often limited to the size of the search radius in the correlation volume, and can not obtain a global field of view. To overcome this problem, we propose the Global Iterative Refinement Module (GIR), which has a detailed structure as shown in Fig.5. Similar to Unet, GIR consists of funnel-shaped encoders and decoders, and the encoding stage mainly uses ConvNextV2 with a large field of view. Unlike the commonly used GRU iteration module, multi-scale large kernel convolution and pyramid pooling module can ensure that GIR obtains sufficient global perception capability.
III-D2 Initial State
The initial states of the optical flow field $(f_{1}^{0},f_{2}^{0})$ and scale change field $f_{3}^{0}$ are given by the initialization network in Sec. III-B, and the initial hidden layer features $h_{0}$ is given by the context encoder.
III-D3 Updating
We iteratively optimize the 3D flow field based on residual update $f^{k+1}=f^{k}+\Delta f$. Before starting the update, we first need to encode the correlation features $F_{scale},F_{of}$, as well as the optical flow field $(f_{1},f_{2})$ and scale change field $f_{3}$, into motion features $F_{motion}$:
$F_{motion}=\textbf{{MotionEncoder}}(F_{scale},F_{of},f_{1},f_{2},f_{3})$ (9)
where MotionEncoder is an encoder composed of a simple two-layer convolution, $F_{motion}\in\mathbb{R}^{H\times W\times 128}$.
Next, as shown in Fig. 5, use the GIR module to iteratively update the hidden features:
$h_{t}=\textbf{{GIR}}(F_{motion},F_{context},h_{t-1})$ (10)
Finally, the residual $\Delta f=(\Delta f_{1},\Delta f_{2},\Delta f_{3})$ for updating is estimated from $h_{t}$ by a simple two-level convolution head:
$\begin{split}(\Delta f_{1},\Delta f_{2})&=\textrm{Conv}(\textrm{ReLU}(\textrm{Conv}(h_{t})))\\ \Delta f_{3}&=\tanh(\textrm{Conv}(\textrm{ReLU}(\textrm{Conv}(h_{t}))))\end{split}$ (11)
Considering that the scale flow requires more detailed estimation, we performed an additional update on $f_{3}$ individually using the GIR network at the end.
III-E Hybrid Training Pipeline
@@
Note that the captured image blocks here are random textures, as we found that using foreground flying objects with clear semantic information (such as complete car, human, and animal contours) does not perform as well as random textures. Specifically, when occlusion occurs, the network becomes overly confident in inferring the motion of the occluded part of the object as planar motion. Random textures can encourage networks to calculate the motion of these flying foreground based on texture matching, and balance the learning weight with occlusion regression.
Calculate Foreground Labels. Before calculating the label for the flight foreground, it is necessary first to determine whether there is an occlusion in the foreground. Because the flight foreground is primarily a random texture, the network cannot make reasonable inferences about the occluded parts. We propose a difference rate $Df$ indicator to measure whether the foreground is retained:
@@
where $N_{q}$ and $N_{r}$ represent the number of foreground flying object pixels in the first and second frames, respectively. In the experiment, we only used flight prospects with $Df$ less than 0.5 to ensure no significant obstructions or great scale changes during the flight.
@@
Optical Flow Loss. We supervise the optical flow results with $L_{1}$ distance between prediction and ground truth, the optical flow loss $\mathcal{L}_{f}$ is defined as:
@@
we set $\gamma=0.8$, $N=6$ in our experiments, the $\bm{f_{gt}}=(f_{1gt},f_{2gt})$ is composed of a hybrid of truth labels from the dataset and the generated $\bm{f_{fg}}$ labels.
@@
where $\bm{\tau}_{gt}$ is also a hybrid of true values $\bm{\tau}$ and generated labels $\bm{\tau}_{fg}$, $\bm{\tau}=Z^{\prime}_{gt}/Z_{gt}$, $Z^{\prime}_{gt}$ and $Z_{gt}$ are the ground truth depths of matching pixels in the second frame and the first frame, we set $\gamma=0.8$, $N=7$ in our experiments.
@@
III-F Convert ScaleFlow++ output to 3D motion
Referring to previous work, we converted the 3D motion $(f_{1},f_{2},f_{3})$ into motion-in-depth and scene flow. Firstly, based on yang’s derivation[20], we can obtain the following approximate equation:
@@ III-F Convert ScaleFlow++ output to 3D motion
where $\bm{Cd}$ is the image coordinate index matrix, normalize scene flow $\overline{\bm{t}}=\bm{K^{-1}}[(f_{3}(\bm{p})-1)\bm{p}+f_{3}(\bm{p})\bm{f}]$. We have presented a detailed derivation process in the Supplementary Material.
TABLE I: Training Details. Pre represents the pre-training stage, and Ft represents the fine-tuning stage. The hybrid training pipeline is used by default during training, and the number of flying foreground additions is 1. Hybrid represents the number of foreground in training
Section Dataset Stage Batch Refine Iteration Train Iteration Size LR Hybrid
Ablation (IV-B) D+ S-18 4 6 10k 320x720 0.000125 1
Motion-in-depth (IV-C) D+T+S+M+K-160 Pre 4 6 30k 320x720 0.00025 1
K-160 Ft 6 6 5k 320x896 0.00005 2
Sintel-test (IV-D) D+T+S+M+K15 Pre 4 6 30k 320x720 0.00025 1
T+S+M Ft 6 6 30k 320x896 0.00005 2
Kitti-test (IV-E) D+T+S+M+K15 Pre 4 6 30k 320x720 0.00025 1
K15 Ft 6 6 5k 320x896 0.00005 2
Generalization (IV-G) ADFactory[59] 4 6 30k 320x768 0.00025 1
@@ IV Experiments
In this section, we introduce the key details of the dataset and experiment, and then further discuss the experimental results. Specifically, the experiment includes the following aspects. Firstly, a comprehensive ablation was conducted on the proposed module. Secondly, test the performance of ScaleFlow++ on mainstream monocular 3D task MID. Thirdly, submit the results of ScaleFlow++ to public testing platform and compared with the most advanced scene flow and optical flow methods currently available. Finally, we qualitatively evaluate the generalization performance of ScaleFlow++ in unseen scenarios.